# Eigenfunction asymptotics in the complex domain for a compact Lie group

### v1

@@ Abstract
Let $(G,\kappa)$ be a compact connected Lie group endowed with a biinvariant Riemannian metric, and let $\tilde{G}$ be the complexification of $G$. We apply Grauert tube techniques to the near-diagonal scaling asymptotics of certain operator kernels, which are defined in terms of the matrix elements of an irreducuble representation drifting to infinity along a ray in weight space. These kernels are the equivariant components of Poisson and Szegő kernels on a fixed sphere bundle in $\tilde{G}$, when the latter is identified with the tangent bundle of $G$ in an appropriate way.
@@ 1 Introduction
is a Fourier integral operator of degree $-(d-1)/2$ with the same canonical relation as $\Pi^{\tau}$. Furthermore, $P^{\tau}$ is closely related to the so-called Poisson-wave operator, which is obtained by holomorphically extending the kernel of the wave operator on $G$ and governs the holomorphic extension of Laplacian eigenfunctions. For this reason, (13) plays a crucial role in the asymptotic study of complexified eigenfunction of compact real-analytic Riemannian manifolds (see also the surveys [Z-2013] and [Z-2017]). On the other hand, (12) is the $\boldsymbol{\lambda}$-th component of $P^{\tau}$; hence it is a natural candidate for the asymptotic study of complexified isotypical eigenfunctions.
In this paper, we shall be concerned with the near-diagonal asymptotics of (11) and (12), when the weight $\boldsymbol{\lambda}$ drifts to infinity along a ray in weight space. In other words, for a fixed non-zero $\boldsymbol{\lambda}\in\mathcal{D}^{G}$, we restrict attention to the ladder of representations $V_{k\,\boldsymbol{\lambda}}$, $k=1,2,\ldots$ (see [GStb]), and consider the asymptotics of the sequence of smoothing kernels $\Pi^{\tau}_{k\,\boldsymbol{\lambda}}$ and $P^{\tau}_{k\,\boldsymbol{\lambda}}$ when $k\rightarrow+\infty$. Analogues of these asymptotics in the line bundle setting were considered in [GsP-2019], [GsP-2020], and [P-2022]. In particular, in spite of the rather different geometric context, the approach in [P-2022] is close to the one in the present work, and the statements bear a formal similarity to those in loc. cit. (with some notational differences).
@@
    $\tilde{G}_{\mathcal{O}}:=\gamma\left(G\times\mathcal{C}(\tilde{\mathcal{O}})\right).$
@@
Let $\jmath^{\tau}:X^{\tau}\hookrightarrow\tilde{G}$ be the inclusion, and set $\alpha^{\tau}:={\jmath^{\tau}}^{*}(\alpha)$; thus $(X^{\tau},\alpha^{\tau})$ is a contact manifold, and $\mathcal{H}_{x}^{\tau}:=\ker(\alpha^{\tau}_{x})$ is the maximal complex subspace of $T_{x}X^{\tau}$ (and has complex dimension $d-1$). Let $\mathcal{R}^{\tau}$ be the Reeb vector field of $(X^{\tau},\alpha^{\tau})$. Then
@@
It follows that NHLC’s at $x\in X^{\tau}$ determine an isomorphism ×$\mathbb{R}\times\mathbb{R}^{2d-2}\cong T_{x}X^{\tau}$, with ×$\mathbb{R}\times\{\mathbf{0}\}$ (and $\{0\}\times\mathbb{R}^{2d-2}$) mapping to, respectively, $\mathrm{span}_{\mathbb{R}}\big{(}\mathcal{R}^{\tau}(x)\big{)}$ and $\mathcal{H}_{x}$. Furthermore, with the identification $\{0\}\times\mathbb{R}^{2d-2}\cong\mathbb{C}^{d-1}$ we obtain an isomorphism of complex vector spaces $\mathbb{C}^{d-1}\cong\mathcal{H}_{x}$.
@@
    Given $x\in\tilde{G}$, let $\mathrm{val}_{x}:\mathfrak{g}\rightarrow T_{x}\tilde{G}$ be the map $\boldsymbol{\xi}\mapsto\boldsymbol{\xi}_{\tilde{G}}(x)$, where $\boldsymbol{\xi}_{\tilde{G}}\in\mathfrak{X}(\tilde{G})$ is the vector field induced by $\boldsymbol{\xi}$ under the action $\tilde{L}$; then $\mathrm{val}_{x}^{t}(\tilde{\kappa}_{x})$ is an Euclidean scalar product on $\mathfrak{g}$.
@@ 2 Notation
$\boldsymbol{\xi}_{\tilde{G}}(x):=\left.\frac{\mathrm{d}}{\mathrm{d}t}\,e^{t\,\boldsymbol{\xi}}\cdot x\right|_{t=0},\quad\boldsymbol{\xi}^{\tilde{G}}(x):=\left.\frac{\mathrm{d}}{\mathrm{d}t}\,x\cdot e^{t\,\boldsymbol{\xi}}\right|_{t=0}\quad(x\in\tilde{G}).$
@@
    $\boldsymbol{\xi}_{\tilde{G}}\in\mathfrak{X}(\tilde{G})$: the vector field induced by $\boldsymbol{\xi}\in\mathfrak{g}$ where $G$ acts on $\tilde{G}$ by left translations (see above);
@@
    $\sigma_{x}:=\mathrm{val}_{x}^{t}(\hat{\kappa}_{x})$ (the Euclidean product on $\mathfrak{g}$ given by pull-back of $\hat{\kappa}$ under vector field evaluation at $x$ see §4.5);
@@ 3 An example
Consider the compact torus $T^{d}=\mathbb{R}^{d}/(2\,\pi\,\mathbb{Z}^{d})$ with the standard metric; in this case $d=r_{T^{d}}$. Let us identify $\mathbb{R}^{d}\cong(\mathbb{R}^{d})^{\vee}$ by the standard Euclidean product. The the unitary dual $\hat{T}^{d}\cong\mathbb{Z}^{d}$, and for each $\boldsymbol{\lambda}\in\mathbb{Z}^{d}$ the subspace $L^{2}(T^{d})_{\boldsymbol{\lambda}}$ is $1$-dimensional. Furthermore, the cone $\mathcal{C}(\mathcal{O})$ is in this case just the positive ray $\mathbb{R}_{+}\cdot\boldsymbol{\lambda}$.
@@
We have $\langle\boldsymbol{\lambda},-\boldsymbol{\omega}\rangle-\|\boldsymbol{\lambda}\|\leq 0$ for any $\boldsymbol{\omega}$ of unit norm, and equality holds if and only if $-\boldsymbol{\omega}=\boldsymbol{\lambda}/\|\boldsymbol{\lambda}\|$; in terms of the moment map $\tilde{\Phi}$ discussed in §4, this is the condition $e^{\imath\,\boldsymbol{\vartheta}+\tau\,\boldsymbol{\omega}}\in{\tilde{\Phi}}^{-1}(\mathbb{R}_{+}\cdot\boldsymbol{\lambda})$.
@@
[…]
@@
[…]
@@
Hence, to leading order,
$\left|\rho^{\tau}_{k\,\boldsymbol{\lambda}}\left(e^{\imath\,\boldsymbol{\vartheta}+\tau\,\boldsymbol{\omega}}\right)\right|^{2}$ $\sim$ $\frac{1}{(2\,\pi)^{d}}\,\left(\frac{k\,\|\boldsymbol{\lambda}\|}{\tau\,\pi}\right)^{\frac{d-1}{2}}\,e^{2\,k\,\tau\,\left[\langle\boldsymbol{\lambda},-\boldsymbol{\omega}\rangle-\|\boldsymbol{\lambda}\|\right]},$
which is also rapidly decreasing unless $-\boldsymbol{\omega}=\boldsymbol{\lambda}/\|\boldsymbol{\lambda}\|$, and in that case it agrees with the first asymptotic expansion in Theorem 1.9.
@@ 4 Preliminaries
In this section, we collect some basic results that will be used in the following proofs. More specific prelimiraries will be given in the subsections devoted to the proofs.
@@ 4.2 The splitting of $T\tilde{G}$
For any $\boldsymbol{\xi}\in\mathfrak{g}$, let $\boldsymbol{\xi}_{\tilde{G}}\in\mathfrak{X}(\tilde{G})$ denote the induced vector field on $\tilde{G}$ under $\tilde{L}$. Then $\boldsymbol{\xi}_{\tilde{G}}$ is Hamiltonian and is tangent to $X^{\tau}$ for every $\tau$. In view of this and (30), by the discussion in §3.2 of [Pao 24] on $\tilde{G}\setminus G$ we have
@@
since both spaces have complex dimension $d-1$, equality holds. ∎
@@ 4.3 The normal bundle to $\tilde{G}_{\mathcal{O}}$ and $X^{\tau}_{\mathcal{O}}$
By (28), the normal bundle $N(X^{\tau}_{\mathcal{O}}/X^{\tau})$ of $X^{\tau}_{\mathcal{O}}$ in $X^{\tau}$ is the restriction to $X^{\tau}$ of the normal bundle $N(\tilde{G}_{\mathcal{O}}/\tilde{G})$ to $\tilde{G}_{\mathcal{O}}$ in $\tilde{G}$. If the stabilizer $T_{\boldsymbol{\lambda}}\leqslant T$ of $\boldsymbol{\lambda}$ has dimension $r_{\boldsymbol{\lambda}}$, then $\mathcal{O}\cong G/T_{\boldsymbol{\lambda}}$ has dimension $d-r_{\boldsymbol{\lambda}}$; hence $\dim\big{(}\mathcal{C}(\mathcal{O})\big{)}=d-r_{\boldsymbol{\lambda}}+1$. By (27), $\dim(\tilde{G}_{\mathcal{O}})=2d-r_{\boldsymbol{\lambda}}+1$; thus $N(\tilde{G}_{\mathcal{O}}/\tilde{G})$ has rank $r_{\boldsymbol{\lambda}}-1$. In the sequel we shall also set $r_{\mathcal{O}}:=r_{\boldsymbol{\lambda}}$ if $\boldsymbol{\lambda}\in\mathcal{O}$.
@@ 4.4 The geodesic flow on $TG$ and $\tilde{G}$
Somewhat more precisely, let $\tilde{\Gamma}^{G}\in\mathfrak{X}(\tilde{G})$ be the vector field on $\tilde{G}$ correlated to $\Gamma^{G}$ by $E$. We shall refer to $\tilde{\Gamma}^{G}$ as the geodesic vector field on $\tilde{G}$. By Remark 4.9,
@@ 4.6.1 The description by Fourier integral operators
As proved by Zelditch, the operator $P^{\tau}$ in (13) is also a Fourier integral operator with complex phase, associated to the same canonical relation as $\Pi^{\tau}$ hence with the same wave front and singular support - and of degree $-(d-1)/2$ (see e.g. [Z-2012], [Z-2014] and [Z-2020], and the discussions in [CR1] and [CR2]). In fact $P^{\tau}=U_{\mathbb{C}}(2\,\imath\,\tau)$, where for general $t\in\mathbb{R}$ one denotes $U_{\mathbb{C}}(t+2\,\imath\,\tau):\mathcal{O}(X^{\tau})\rightarrow\mathcal{O}(X^{\tau})$ the complexified Poisson-wave operators obtained (roughly speaking) by holomorphically extending the operator kernel of the wave operator. Zelditch showed that complexified Poisson-wave operators may described in terms of dynamical Toeplitz operators; one can then conclude that $P^{\tau}$ is a Toeplitz operator of degree $-(d-1)/2$, given by the compression by $\Pi^{\tau}$ of a pseudodifferential operator $Q^{\tau}$ of the same degree. To leading order, $Q^{\tau}$ has the form $(\pi\,\tau)^{\frac{d-1}{2}}\cdot D_{\sqrt{\rho}}^{-(d-1)/2}$ (cfr. the discussion in §5 of [GvP-2024]). It follows that the operator kernel of $P^{\tau}$ has the form
@@ 5 Proofs of the Theorems
[…]
@@
Let $R_{+}\subset R$ denote the collection of positive roots of the pair $(\mathfrak{g},\mathfrak{t})=(\hat{\mathfrak{g}},\hat{\mathfrak{t}})$, and set

### v2

@@ Abstract
Let $(G,\kappa)$ be a compact connected Lie group endowed with a biinvariant Riemannian metric, and let $\tilde{G}$ be the complexification of $G$. We apply Grauert tube techniques to the near-diagonal scaling asymptotics of certain operator kernels, which are defined in terms of the matrix elements of an irreducible representation drifting to infinity along a ray in weight space. These kernels are the equivariant components of Poisson and Szegő kernels on a fixed sphere bundle in $\tilde{G}$, when the latter is identified with the tangent bundle of $G$ in an appropriate way.
@@ 1 Introduction
is a Fourier integral operator of degree (d1)/2-(d-1)/2 with the same canonical relation as $\Pi^{\tau}$. Furthermore, $P^{\tau}$ is closely related to the so-called Poisson-wave operator, which is obtained by holomorphically extending the kernel of the wave operator on $G$ and governs the holomorphic extension of Laplacian eigenfunctions. For this reason, (13) plays a crucial role in the asymptotic study of complexified eigenfunction of compact real-analytic Riemannian manifolds (see also the surveys [Z-2013] and [Z-2017]). On the other hand, (12) is the $\boldsymbol{\lambda}$-th component of $P^{\tau}$; hence it is a natural candidate for the asymptotic study of complexified isotypical eigenfunctions.
In this paper, we shall be concerned with the near-diagonal asymptotics of (11) and (12), when the weight $\boldsymbol{\lambda}$ drifts to infinity along a ray in weight space. In other words, for a fixed non-zero $\boldsymbol{\lambda}\in\mathcal{D}^{G}$, we restrict attention to the ladder of representations $V_{k\,\boldsymbol{\lambda}}$, $k=1,2,\ldots$ (see [GStb]), and consider the asymptotics of the sequence of smoothing kernels $\Pi^{\tau}_{k\,\boldsymbol{\lambda}}$ and $P^{\tau}_{k\,\boldsymbol{\lambda}}$ when $k\rightarrow+\infty$. Analogues of these asymptotics in the line bundle setting were considered in [GsP-2019], [GsP-2020], and [P-2022]. In particular, in spite of the rather different geometric context, the approach in [P-2022] is close to the one in the present work, and the statements bear a formal similarity to those in loc. cit. (see the closing remark of this introduction).
@@
[…]
@@
Let $\jmath^{\tau}:X^{\tau}\hookrightarrow\tilde{G}$ be the inclusion, and set $\alpha^{\tau}:={\jmath^{\tau}}^{*}(\alpha)$; thus $(X^{\tau},\alpha^{\tau})$ is a contact manifold, and $\mathcal{H}_{x}^{\tau}:=\ker(\alpha^{\tau}_{x})$ is the maximal complex subspace of $T_{x}X^{\tau}$ (and has complex dimension d1d-1). Let $\mathcal{R}^{\tau}$ be the Reeb vector field of $(X^{\tau},\alpha^{\tau})$. Then
@@
It follows that NHLC’s at $x\in X^{\tau}$ determine an isomorphism $\mathbb{R}\times\mathbb{R}^{2d-2}\cong T_{x}X^{\tau}$, with $\mathbb{R}\times\{\mathbf{0}\}$ (and $\{0\}\times\mathbb{R}^{2d-2}$) mapping to, respectively, $\mathrm{span}_{\mathbb{R}}\big{(}\mathcal{R}^{\tau}(x)\big{)}$ and $\mathcal{H}_{x}$. Furthermore, with the identification $\{0\}\times\mathbb{R}^{2d-2}\cong\mathbb{C}^{d-1}$ we obtain an isomorphism of complex vector spaces $\mathbb{C}^{d-1}\cong\mathcal{H}_{x}$.
@@
[…]
@@ 1.1 Applications
Let us close this introduction with a remark on a slight difference in the representation-theoretic approach between [P-2022], where similar asymptotics were considered in the line bundle setting, and the present paper. Given a dominant weight $\boldsymbol{\lambda}$, let us set $\boldsymbol{\nu}:=\boldsymbol{\nu}_{\boldsymbol{\lambda}}=\boldsymbol{\lambda}+\boldsymbol{\delta}$, where $\boldsymbol{\delta}$ is the half-sum of the positive roots - see (57) below; for a fixed dominant weight $\boldsymbol{\lambda}$, the equivariant asymptotics in [P-2022] are those associated to the sequence of weights $\boldsymbol{\lambda}_{k}$ such that $\boldsymbol{\nu}_{\boldsymbol{\lambda}_{k}}=k\,\boldsymbol{\nu}$.
In this paper, given a fixed regular dominant weight $\boldsymbol{\lambda}$, we consider the asymptotics associated to the representations $k\,\boldsymbol{\lambda}$. This implies that the role played by $\mathcal{O}_{\boldsymbol{\nu}}$ in [P-2022] is played here by $\mathcal{O}_{\boldsymbol{\lambda}}$.
This approach is perhaps more natural, and makes the character computations based on the Kirillov character formula only slightly more involved. The same arguments and change in approach could of course be applied to the setting of [P-2022].
@@ 2 Notation
[…]
@@
[…]
@@
    $\sigma_{x}:=\mathrm{val}_{x}^{t}(\hat{\kappa}_{x})$ (the Euclidean product on $\mathfrak{g}$ given by pull-back of $\hat{\kappa}$ under vector field evaluation at $x$ - see §4.5);
@@ 3 An example
Consider the compact torus $T^{d}=\mathbb{R}^{d}/(2\,\pi\,\mathbb{Z}^{d})$ with the standard metric (cfr. §2 of [P-2024]); in this case $d=r_{T^{d}}$. Let us identify $\mathbb{R}^{d}\cong(\mathbb{R}^{d})^{\vee}$ by the standard Euclidean product. The the unitary dual $\hat{T}^{d}\cong\mathbb{Z}^{d}$, and for each $\boldsymbol{\lambda}\in\mathbb{Z}^{d}$ the subspace $L^{2}(T^{d})_{\boldsymbol{\lambda}}$ is $1$-dimensional. Furthermore, the cone $\mathcal{C}(\mathcal{O})$ is in this case just the positive ray $\mathbb{R}_{+}\cdot\boldsymbol{\lambda}$.
@@
We have $\langle\boldsymbol{\lambda},-\boldsymbol{\omega}\rangle-\|\boldsymbol{\lambda}\|\leq 0$ for any $\boldsymbol{\omega}$ of unit norm, and equality holds if and only if $\boldsymbol{\omega}=\boldsymbol{\lambda}/\|\boldsymbol{\lambda}\|$; in terms of the moment map $\tilde{\Phi}$ discussed in §4, this is the condition $e^{\imath\,\boldsymbol{\vartheta}+\tau\,\boldsymbol{\omega}}\in{\tilde{\Phi}}^{-1}(\mathbb{R}_{+}\cdot\boldsymbol{\lambda})$.
@@
$\Psi_{\boldsymbol{\lambda}}(\boldsymbol{\omega})=\imath\,2\,\tau\,\left[\|\boldsymbol{\lambda}\|-\langle\boldsymbol{\lambda},\boldsymbol{\omega}\rangle\right].$
@@
$\sqrt{\det\left(\frac{1}{2\,\pi\,\imath}\,2\,\imath\,\tau\,\|\boldsymbol{\lambda}\|\,I_{d-1}\right)}=\sqrt{\left(\frac{\tau\,\|\boldsymbol{\lambda}\|}{\pi}\right)^{d-1}}=\left(\frac{\tau\,\|\boldsymbol{\lambda}\|}{\pi}\right)^{\frac{d-1}{2}}.$
@@
Therefore, to leading order,
$\left|\rho^{\tau}_{k\,\boldsymbol{\lambda}}\left(e^{\imath\,\boldsymbol{\vartheta}+\tau\,\boldsymbol{\omega}}\right)\right|^{2}$ $\sim$ $\frac{1}{(2\,\pi)^{d}}\,\left(\frac{k\,\|\boldsymbol{\lambda}\|}{\tau\,\pi}\right)^{\frac{d-1}{2}}\,e^{2\,k\,\tau\,\left[\langle\boldsymbol{\lambda},-\boldsymbol{\omega}\rangle-\|\boldsymbol{\lambda}\|\right]},$ (25)
which is also rapidly decreasing unless $\boldsymbol{\omega}=\boldsymbol{\lambda}/\|\boldsymbol{\lambda}\|$, and in that case it agrees with the first asymptotic expansion in Theorem 1.9.
Let us consider a rescaled displacement in a normal direction to $Z^{\tau}$, which amounts in this case to replacing (to first order) $\tau\,\boldsymbol{\omega}=\tau\,(-\boldsymbol{\lambda}/\|\boldsymbol{\lambda}\|)$ by $\tau\,\boldsymbol{\omega}=\tau(-\boldsymbol{\lambda}/\|\boldsymbol{\lambda}\|)+\mathbf{n}/\sqrt{k}$, where $\mathbf{n}\in\boldsymbol{\lambda}^{\perp}$ is fixed. In the previous general notation, we consider the asymptotics at a diagonal pair $(x+\mathbf{n}/\sqrt{k},x+\mathbf{n}/\sqrt{k})$, with $x\in X^{\tau}_{\mathcal{O}}$ and $\mathbf{n}$ normal to $X^{\tau}_{\mathcal{O}}$. Since $\boldsymbol{\omega}\in S^{d-1}$,
$\boldsymbol{\omega}=\sqrt{1-\frac{1}{\tau^{2}\,k}\|\mathbf{n}\|^{2}}\,\left(-\frac{\boldsymbol{\lambda}}{\|\boldsymbol{\lambda}\|}\right)+\frac{1}{\tau\,\sqrt{k}}\,\mathbf{n}.$
Hence, the exponent in (25) is
$2\,k\,\tau\,\left[\langle\boldsymbol{\lambda},-\boldsymbol{\omega}\rangle-\|\boldsymbol{\lambda}\|\right]=2\,k\,\tau\,\|\boldsymbol{\lambda}\|\,\left[\sqrt{1-\frac{1}{\tau^{2}\,k}\|\mathbf{n}\|^{2}}-1\right]$
$=$ $2\,k\,\tau\,\|\boldsymbol{\lambda}\|\,\left[-\frac{1}{2\,\tau^{2}\,k}\|\mathbf{n}\|^{2}+R_{2}\left(\frac{\|\mathbf{n}\|^{2}}{k}\right)\right]=\frac{\|\boldsymbol{\lambda}\|}{\tau}\,\left[-\|\mathbf{n}\|^{2}+O\left(k^{-1}\right)\right]$
$=$ $\frac{\|\boldsymbol{\lambda}\|}{\tau}\,\left[-2\,\|\mathbf{n}\|_{\tilde{\kappa}}^{2}+O\left(k^{-1}\right)\right],$
where $\|\cdot\|^{2}_{\tilde{\kappa}}=\|\cdot\|^{2}/2$ in the notation of the introduction. The same argument applies for, say, $\|\mathbf{n}\|=O\left(k^{1/6}\right)$, with a remainder term $O\left(k^{-1/3}\right)$.
@@ 4 Preliminaries
In this section, we collect some basic results that will be used in the following proofs. More specific preliminaries will be given in the subsections devoted to the proofs.
@@ 4.2 The splitting of $T\tilde{G}$
[…]
@@
since both spaces have complex dimension d1d-1, equality holds. ∎
@@ 4.3 The normal bundle to $\tilde{G}_{\mathcal{O}}$ and $X^{\tau}_{\mathcal{O}}$
[…]
@@ 4.4 The geodesic flow on $TG$ and $\tilde{G}$
[…]
@@ 4.6.1 The description by Fourier integral operators
As proved by Zelditch, the operator $P^{\tau}$ in (13) is also a Fourier integral operator with complex phase, associated to the same canonical relation as $\Pi^{\tau}$ - hence with the same wave front and singular support - and of degree (d1)/2-(d-1)/2 (see e.g. [Z-2012], [Z-2014] and [Z-2020], and the discussions in [CR1] and [CR2]). In fact $P^{\tau}=U_{\mathbb{C}}(2\,\imath\,\tau)$, where for general $t\in\mathbb{R}$ one denotes $U_{\mathbb{C}}(t+2\,\imath\,\tau):\mathcal{O}(X^{\tau})\rightarrow\mathcal{O}(X^{\tau})$ the complexified Poisson-wave operators obtained (roughly speaking) by holomorphically extending the operator kernel of the wave operator. Zelditch showed that complexified Poisson-wave operators may described in terms of dynamical Toeplitz operators; one can then conclude that $P^{\tau}$ is a Toeplitz operator of degree (d1)/2-(d-1)/2, given by the compression by $\Pi^{\tau}$ of a pseudodifferential operator $Q^{\tau}$ of the same degree. To leading order, $Q^{\tau}$ has the form $(\pi\,\tau)^{\frac{d-1}{2}}\cdot D_{\sqrt{\rho}}^{-(d-1)/2}$ (cfr. the discussion in §5 of [GvP-2024]). It follows that the operator kernel of $P^{\tau}$ has the form
@@ 5 Proofs of the Theorems
Let us consider $U_{2}$. Let $L_{\boldsymbol{\lambda}}=(\boldsymbol{\lambda},\,2\,\boldsymbol{\lambda},\,\ldots)$ denote the ladder of irreducible representations generated by $\boldsymbol{\lambda}$ [GStb]. Let us consider the corresponding subspaces
@@
[…]